# Light Curve Classification with DistClassiPy: a new distance-based classifier

### v1

@@ 5.1 Classification Problems
Figure 6: The canonical hierarchy tree for the ten classes included in our work. This is a physical classification, based on (Eyer and Mowlavi, 2008). Note: the grey dot represents the objects which belong to none of the other classes, which are not represented in our dataset.
Our analysis will focus on the more comprehensive multi-class classification, where we look at four classes of Cepheid-like pulsating variables: Cepheid (CEP), RR Lyrae Type a/b (RR), RR Lyrae Type c (RRc), and Delta Scuti (DSCT). We discuss in detail these results in section 6. But we also report the performance of our method on a specific binary case and One vs. Rest case for comparison, keeping in mind that the results obtained when choosing one or two classes may differ significantly for any other class or class combination. In the One vs. Rest classification, we isolate Eclipsing Algols (EA) from all other classes. In the binary classification, we separate two classes of rotating variables (RS Canum Venaticorum, RSCVN, and BY Draconis, BYDra). This is a particularly challenging classification because while physically different, both classes exhibit photometrically similar behaviors, with variability at the $\sim 0.1-0.2$ mag level, and with varying phase, magnitude slope, amplitude, and period (Bopp, 1980).
@@ 5.2 DistClassiPy Classification Algorithm
We use a custom algorithm, which we have named DistClassiPy, to classify the light curves. Our method draws inspiration from the $k$-Nearest Neighbours ($k-$NN; (Cover and Hart, 1967; Fix and Hodges, 1989)) algorithm. In addition to the classification that $k-$NN provides, DistClassiPy also offers a quantification of the uncertainty in the prediction. We have implemented this in Python using the scikit-learn API (Buitinck et al., 2013).
@@ 5.2.1 Training
We compute the median and standard deviation for each feature per class, where we use the median set as a representative for each class [^10] The median and standard deviation can also be replaced by other statistical measures of central tendency and dispersion.. Using the median here also makes our calculations more resilient to outliers. An illustration of the median values for four representative features can be found in Table 2.
@@
Class $r_{\text{period}}$ $r_{\text{amplitude}}$ $r_{\text{R21}}$ $r_{\text{skew}}$
CEP 3.86 0.25 0.27 -0.16
RR 0.56 0.35 0.45 -0.52
RRc 0.33 0.20 0.15 0.01
DSCT 0.09 0.08 0.17 -0.14
@@
Figure 7: How DistClassiPy performance changes for the multi-class classification problem with the number of features used. Here we show our final step in dimensionality reduction, the Sequential Feature Selection (SFS) process, for two sample distance metrics: Clark (Figure 7a) and Canberra (Figure 7b). The performance of the classifier, measured as the mean $F_{1}$ score in a 5-fold cross validation run, is shown as a function of the number of features included in the classification (with the standard deviation shown as a filled area). The SFS algorithm iteratively adds one feature at a time from $n$=1 to $n$=31 such that the $F_{1}$ score is maximized at each step. We then select a final set of $n_{\mathrm{final}}$ features such that $n_{\mathrm{final}}$ is the lowest number of features whose $F_{1}$ score is within $1\sigma$ of the maximum $F_{1}$ score (denoted by a red vertical line in each panel).
@@ 5.3 Classification Confidence
First, as is standard in the ML community, we assess the stability of our classification results with cross-validation (Kohavi, 1995): we split our dataset into 5-folds, use four folds as our training set and one fold as our testing set, and average the five $F_{1}$ scores to obtain the final score. This assesses the confidence of the result for a given dataset and the influence of data outliers, but does not provide a direct measure of the uncertainty associated with the use of a specific distance metric.
Therefore, we developed three different confidence parameters for use with the classifier. The goal was to develop a distance-based metric that measures uncertainty (or confidence) in the classification method, not just the data. We tested the performance of the following three distance-derived confidence measures:
Figure 9: The top features for the multi-class classification task across all metrics. This plot shows 15 features that were included in the final set of $n_{\mathrm{final}}$ features for $\geq 3$ metrics. All but one metric (the Maryland Bridge metric) identified Period_band_r as important.
@@ 5.4 Sequential Feature Selection and Classification
We have already reduced our feature space from 112 to 31 features. To increase the robustness of our method and decrease the computational cost, we further reduce our dimensionality by selecting the most effective classification features for each class and each problem. This process is computationally expensive (as described below) and the final set of features to be used depends on the metric used and the classification task. However, this computation can be performed once for each problem, as we did for the three classifications we explored in our work. Thereafter, the user of our algorithm can choose the features that best match their research based on their scientific interest and object sample, with a net computational gain. Here, we describe our strategy, and the detailed results are discussed in section 6.3.
@@
In Figure 9 we show the number of distance metrics (out of our 18 metrics) for which a given feature is selected. The 15 features shown were important for $\geq 3$ metrics and are considered as a “super-set” of important features.
@@ 6.1 DistClassiPy classification performance
[…]
Finally, we test our final classifiers on the hidden set introduced in section 3.1, which consists of $500$ objects, entirely unseen in training. For each classification, we test the best model on a hidden dataset, first rebalanced and then unbalanced, where each class has a representation similar to the fraction of objects in the initial dataset. We confirmed the model performance for all three classification tasks, and found the results to be mostly within $1\sigma$ expectations from our cross-validation results for the rebalanced dataset. With the best distance metric selected on the cross-validated data, for each classification task the performance on the hidden set is $F_{1}=93\%$ with the Clark metric on the multi-class classification (1881 objects), $F_{1}=68\%$ with the Euclidean metric on the binary classification (1000 objects) and $F_{1}=92\%$ with the Kulczynski metric for the One vs. Rest classification (976 objects; section 5). For the unbalanced dataset, the performances change slightly. In the multi-class classification, the unbalanced dataset contains 494 RR, 253 DSCT, 210 RRc, 24 CEP. The performance of the Canberra-based model improves slightly from $F_{1}=91\%$ to $F_{1}=93\%$, while the performance of the Clark-based model drops to $F_{1}=86\%$ from $92\%$. This is primarily due to confusion between the RR and RRc classes, which are the most common labels. However, for comparison, the RFC performance drops significantly more, from $F_{1}=92\%$ to $F_{1}=76\%$. In the imbalanced One vs. Rest classification, the performance drops by a few points, from $F_{1}=93\%$ (with Euclidean distance) to $F_{1}=90\%$. Similarly, the RFC performance drops from $F_{1}=95\%$ to $F_{1}=89\%$. The binary classification is naturally balanced with 500 BYDra and 480 RSCVN.
@@
The complete set of results is available online as a spreadsheet. [^12] Hidden Set Results URL. The confusion matrices for the multi-class classification of the hidden set are displayed in Figure 13 for the Clark and Canberra metric.
@@ 6.2 Computational Requirements
To understand how the computational time scales with dataset size and feature dimensions, we monitored the total time taken for training and prediction for different numbers of data points and features using synthetic data generated with scikit-learn - the specifications for our hardware and software are outlined in C. We compared the timings of DistClassiPy using the metrics Canberra and Clark. It’s worth noting that the Canberra metric is implemented in the scipy.spatial.distance package, and we thus expect it to be computationally optimized. On the other hand, the Clark metric is not available within Scipy (to our knowledge) and we coded it ourselves: we expect it may be sub-optimal in terms of computational costs. The timings of an RFC with 100 trees is also tracked. We ran each method five times, and report the mean here. Our results are illustrated in Figure 14 and Figure 15.
@@
Figure 14: Top: A visual comparison of computational times between our method, DistClassiPy and RFC as a function of the dataset size (number of objects; $y$-axis) and number of features ($x$-axis). The area of each circle is proportional to the computational time required for training and generating the predictions. We see that DistClassiPy is always faster than RFC (with 100 trees), and the difference is even more prominent in larger datasets. Bottom: A comparison of computational times (in seconds) between DistClassiPy and RFC as a function of variable dataset size and number of features. We see that DistClassiPy is always at least an order of magnitude faster than RFC (with 100 trees), even when using custom metrics (e.g. Clark) which are not in-built in SciPy (like, e.g., Canberra).
Figure 15: The computational time of DistClassiPy using the Canberra and Clark metrics (top and bottom rows respectively) as a function of sample size using 100 features (left plots) and of number of features using 1000 objects (right plots).
We find that DistClassiPy is remarkably scalable and shows minimal computation times even for large and high-dimensional datasets. Even when dealing with a dataset size of 10,000 objects and 500 features, the computation times for DistClassiPy remain as low as $0.0730\pm 0.0061$ seconds (Canberra) and $0.1558\pm 0.0061$ seconds (Clark) compared to $14.4565\pm 0.3204$ seconds for RFC. These make DistClassiPy very suitable for current and future astrophysical surveys, with increasing sample sizes (e.g., the Vera C. Rubin LSST is expected to characterize 17B stars, many of which are variable to the depth limit of the survey, and 48B astrophysical objects altogether).
Finally, DistClassiPy and RFC utilize a comparable amount of RAM (the distance metric classifier utilizes $359.2\pm 20.3$ MiB compared to RFC’s $380.2\pm 37.6$ MiB, for the dataset having 10,000 objects and 500 features) during training, as per our tests using the Python package memory_profiler.
@@
We now look at other assets that DistClassiPy can offer us to help understand classification - confidence and robustness.
@@ 6.3 Feature Importance
We find that whether a feature is important or not depends not only on the classification task, but also the distance metric being used. In other words, each metric has features with which it works best, and often, these features differ for different metrics. We show in Figure 16 how the important features change for the multi-class classification task for three metrics - Canberra, Clark, and Soergel.
Figure 16: How feature importance changes for three different distance metrics (Canberra, Clark, and Soergel) in the multi-class classification. Each of these metrics has a similar performance, but there’s only one feature that’s common among all of them: Period_band_r - the $r$-band period.
@@ 6.4 Confidence of the distance metric classification
An interesting point to note is that, unlike the probabilistic outputs from something like RFC, our distance metric classifier with $c_{d}$ does not assume a fixed confidence level shared across all classes. The distance metric classifier doesn’t impose a strict boundary on how far away a test object can be, and we suspect this may be useful for dealing with outliers and anomalies. However, this is not something we will explore in this work, and we end our discussion on confidences.
@@ 6.5 Robustness of the distance metric classification
We are interested in determining the robustness of our classifier - how well does our classifier perform with different types of data? We split the test set into batches, splitting the data by a given feature’s distribution into four quantiles, ensuring that each quantile contains an equal number of objects. Then, we evaluate the accuracy for each quantile separately, only using the final set of features from the SFS (section 4). Since this accuracy is now a measure of robustness, we call it the robustness score. This method is applied to all features from our “super-set” (as derived in Figure 9). For the multi-class classification task, robustness scores for the Clark and Canberra metric have been visualized as a heatmap in Figure 18. We note that, although all features in the “super-set” of the 15 best features are not used by the classifier itself, we still use them to split our quantiles, before dropping the unused feature based on the distance metric. This helps us compare the robustness for different metrics.
@@
Figure 18: Robustness scores for the multi-class classification task, with the distance metric classifier using the Clark (Figure 18a) and Canberra metrics (Figure 18b). The robustness score is calculated by splitting our test set into quantiles, and then evaluating the classification accuracy for each quantile. Note that, although each quantile is used to split the dataset, only a subset of these features are actually used in the classification (boldfaced). We still include all features from the super-set of features for the quantile split, to better compare performance for different metrics. For e.g., although the Canberra metric (b) performs better in general, Clark metric (a) performs better with objects having larger periods (quantile 4, top row).
@@ 7 Conclusions
Time series classification is a non-trivial problem in machine learning. In ground-based astronomy, light curve classification is even more challenging due to the sparsity and heteroscedasticity of the data. We developed a distance metric classifier  DistClassiPy that uses distance metrics computed on features derived from light curves. Tested on multiple classification tasks, our classifier produces results comparable to other state-of-the-art feature-based methods like Random Forest Classifiers (RFCs), with $F_{1}=92\%$ and $F_{1}=94\%$ on a multi-class and a One vs. Rest classification respectively. We achieve a performance of $F_{1}=68\%$ on a particularly challenging binary classification that attempts to separate RSCVN and BYDra, two star types that exhibit photometrically similar and highly variable behaviors (Bopp, 1980), and for which, too, the RFC performance is comparable $F_{1}=70\%$. This cross-validated performance was confirmed by running the model on a “hidden” dataset completely unseen in training and that, unlike the original dataset, is not rebalanced to have an even representation of all classes. For all three classification tasks, we confirm the performance obtained in cross-validation testing: $F_{1}=93\%,92\%,$ and $68\%$ respectively for the multi-class, One vs. Rest, and binary classification.
At the same time, DistClassiPy is faster, more interpretable, and suited to be fine tuned to the specific science goals with additional performance enhancements and computational advantages. In our tests, it performed 50×50$\times$50 × faster than RFC on small datasets (10 features by 100 objects), and over 150×150$\times$150 × faster on larger datasets (500 features by 10,000 objects) demonstrating impressive scalability properties. The dynamism in the distance metrics allows for a choice suited to the available features and the test particle’s characteristics within the feature space. Finally, because the features are a result of domain-sensitive dimensionality reduction, the user can leverage relevant knowledge in the context of the problem.

### v2

@@ 5.1 Classification Problems
Figure 6: The canonical hierarchy tree for the ten classes included in our work. This is a physical classification, based on Eyer and Mowlavi (2008). Note: the gray dot represents the objects which belong to none of the other classes, which are not represented in our dataset.
Our analysis will focus on the more comprehensive multi-class classification, where we look at four classes of Cepheid-like pulsating variables: Cepheid (CEP), RR Lyrae Type ab (RR), RR Lyrae Type c (RRc), and Delta Scuti (DSCT). We discuss in detail these results in section 6. But we also report the performance of our method on a specific binary case and One vs. Rest case for comparison, keeping in mind that the results obtained when choosing one or two classes may differ significantly for any other class or class combination. In the One vs. Rest classification, we isolate Eclipsing Algols (EA) from all other classes. In the binary classification, we separate two classes of rotating variables (RS Canum Venaticorum, RSCVN, and BY Draconis, BYDra). This is a particularly challenging classification because while physically different, both classes exhibit photometrically similar behaviors, with variability at the $\sim 0.1-0.2$ mag level, and with varying phase, magnitude slope, amplitude, and period (Bopp, 1980).
@@ 5.2 DistClassiPy Classification Algorithm
We use a custom algorithm, which we have named DistClassiPy, to classify the light curves. Our method draws inspiration from the $k$-Nearest Neighbors ($k-$NN; (Cover and Hart, 1967; Fix and Hodges, 1989)) algorithm. In addition to the classification that $k-$NN provides, DistClassiPy also offers a quantification of the uncertainty in the prediction. We have implemented this in Python using the scikit-learn API (Buitinck et al., 2013).
@@ 5.2.1 Training
Class $r_{\text{period}}$ $r_{\text{amplitude}}$ $r_{\text{R21}}$ $r_{\text{skew}}$
CEP 3.86 0.25 0.27 -0.16
RR 0.56 0.35 0.45 -0.52
RRc 0.33 0.20 0.15 0.01
DSCT 0.09 0.08 0.17 -0.14
We compute the median and standard deviation for each feature per class, where we use the median set as a representative for each class. [^12] The median and standard deviation can also be replaced by other statistical measures of central tendency and dispersion. Using the median here also makes our calculations more resilient to outliers. An illustration of the median values for four representative features can be found in Table 2. The distance metric is not selected in the training step.
@@ 5.3 Classification Confidence
First, as is standard in the machine learning community, we assess the stability of our classification results with cross-validation (Kohavi, 1995): we split our dataset into 5-folds, use four folds as our training set and one fold as our testing set, and average the five $F_{1}$ scores to obtain the final score. This assesses the confidence of the result for a given dataset and the influence of data outliers, but does not provide a direct measure of the uncertainty associated with the use of a specific distance metric.
Therefore, we additionally developed three confidence parameters for use with the classifier. Our goal was to create a distance-based measure for uncertainty (or confidence) in the classification method, not just the data. We tested the performance of the following three distance-derived confidence measures:
Figure 9: The top features for the multi-class classification task across all metrics. This plot shows 15 features that were included in the final set of $n_{\mathrm{final}}$ features for $\geq$3 metrics. All but one metric (the Maryland Bridge metric) identified Period_band_r as important.
@@ 5.4 Sequential Feature Selection and Classification
We have already reduced our feature space from 114 to 31 features. To increase the robustness of our method and decrease the computational cost, we further reduce our dimensionality by selecting the most effective classification features for each class and each problem. This process is computationally expensive (as described below) and the final set of features to be used depends on the metric used and the classification task. However, this computation can be performed once for each problem, as we did for the three classifications we explored in our work. Thereafter, the user of our algorithm can choose the features that best match their research based on their scientific interest and object sample, with a net computational gain. Here, we describe our strategy, and the detailed results are discussed in section 6.3.
@@
In Figure 9 we show the number of distance metrics (out of our 18 metrics) for which a given feature is selected. The 15 features shown were important for $\geq$3 metrics and are considered as a “super-set” of important features. We emphasize that the user should select the most effective distance metric and subset of features, performing a feature selection such as the one described above, specifically for their data and classification problem, as the classification goal (e.g., purity vs completeness in selecting a specific class of objects, vs separating two specific subclasses, etc.), as well as data quality (including sparsity, signal-to-noise ratio, noise property, number of bands, etc.) will influence the selection of features and ultimate performance of the model.
@@ 6.1 DistClassiPy classification performance
The best cross-validation result of our distance metric classifier (DMC) from DistClassiPy across all distance metrics and feature subsets for each classification task are as follows: for the Multi-class Classification, $\text{DMC}=92\%\ \text{(}\text{RFC}=92\%$); for the One vs. Rest, $\text{DMC}=94\%\ \text{(}\text{RFC}=95\%$; for the Binary classification of rotating stars, $\text{DMC}=68\%\ \text{(}\text{RFC}=70\%$. A comprehensive report of the performance of our model, including the distribution of performance values across different distance metric and feature selection choices, is presented in Table 3. The full spreadsheet is available online. [^13] Cross-Val Results URL Confusion matrices for the multi-class classification task for the distance metric classifier (using the top two distance metrics: Clark and Canberra) and RFC are shown in Figure 12. From these, we see that performance is mostly very similar. However, for some classes like RR and RRc, the Canberra distance classifier outperforms the RFC.
Finally, we test our final classifiers on the hidden set introduced in section 3.1, which consists of $500$ objects, entirely unseen in training. For each classification, we test the best model (best distance metric and features subset) on a hidden dataset, first rebalanced and then unbalanced, where each class has a representation similar to the fraction of objects in the initial dataset. We confirmed the model performance for all three classification tasks, and found the results to be mostly within $1\sigma$ expectations from our cross-validation results for the rebalanced dataset. With the best distance metric selected on the cross-validated data, for each classification task the performance on the hidden set is $F_{1}=93\%$ with the Clark metric on the multi-class classification (1881 objects), $F_{1}=68\%$ with the Euclidean metric on the binary classification (1000 objects) and $F_{1}=92\%$ with the Kulczynski metric for the One vs. Rest classification (976 objects; section 5). For the unbalanced dataset, the performances change slightly. In the multi-class classification, the unbalanced dataset contains 494 RR, 253 DSCT, 210 RRc, 24 CEP. The performance of the Canberra-based model improves slightly from $F_{1}=91\%$ to $F_{1}=93\%$, while the performance of the Clark-based model drops to $F_{1}=86\%$ from 92%. This is primarily due to confusion between the RR and RRc classes, which are the most common labels. However, for comparison, the RFC performance drops significantly more, from $F_{1}=92\%$ to $F_{1}=76\%$. In the imbalanced One vs. Rest classification, the performance drops by a few points, from $F_{1}=93\%$ (with Euclidean distance) to $F_{1}=90\%$. Similarly, the RFC performance drops from $F_{1}=95\%$ to $F_{1}=89\%$. The binary classification is naturally balanced with 500 BYDra and 480 RSCVN.
@@
The complete set of results is available online as a spreadsheet.141414[Hidden Set Results URL]. The confusion matrices for the multi-class classification of the hidden set are displayed in Figure 13 for the Clark and Canberra metric.
@@ 6.2 Computational Requirements
To understand how the computational time scales with dataset size and feature dimensions, we monitored the total time taken for training and prediction for different numbers of data points and features using synthetic data generated with scikit-learn — the specifications for our hardware and software are outlined in C. We compared the timings of DistClassiPy using the metrics Canberra and Clark. It is worth noting that the Canberra metric is implemented in the scipy.spatial.distance package, and we thus expect it to be computationally optimized. On the other hand, the Clark metric is not available within Scipy (to our knowledge) and we coded it ourselves: we expect it may be sub-optimal in terms of computational costs. The timings of an RFC with 100 trees is also tracked. We ran each method five times, and report the mean here. Our results are illustrated in Figure 14 and Figure 15.
@@
Figure 14: Top: A visual comparison of computational times between our method, DistClassiPy and RFC as a function of the dataset size (number of objects; $y$-axis) and number of features ($x$-axis). For all plots, the computational costs of fitting the model on the training data and predicting on the testing data (not including preprocessing) are evaluated on the entire dataset five times, and the results shown are the average over these five iterations. The area of each circle is proportional to the computational time required for training and generating the predictions. We see that DistClassiPy is always faster than RFC (with 100 trees), and the difference is even more prominent in larger datasets. Bottom: A comparison of computational times (in seconds) between DistClassiPy and RFC as a function of variable dataset size and number of features. We see that DistClassiPy is always at least an order of magnitude faster than RFC (with 100 trees), even when using custom metrics (e.g. Clark) which are not in-built in SciPy (like, e.g., Canberra).
Figure 15: The computational time of DistClassiPy using the Canberra and Clark metrics (left and right columns respectively) as a function of sample size using 100 features (top plots) and of number of features using 1000 objects (bottom plots).
We find that DistClassiPy is remarkably scalable and shows minimal computation times even for large and high-dimensional datasets. Even when dealing with a dataset size of 10,000 objects and 500 features, the computation times for DistClassiPy remain as low as $0.0677\pm 0.0016$ seconds (Canberra) and $0.1538\pm 0.0023$ seconds (Clark) compared to $14.3563\pm 0.4978$ seconds for RFC. These make DistClassiPy very suitable for current and future astrophysical surveys, with increasing sample sizes (e.g., the Vera C. Rubin LSST is expected to characterize 17B stars, many of which are variable to the depth limit of the survey, and 48B astrophysical objects altogether).
Finally, DistClassiPy and RFC utilize a comparable amount of RAM (the distance metric classifier utilized $574.5\pm 4.8$ MiB compared to RFC’s $568.4\pm 16.3$ MiB on our machine. [^15] Note: These numbers are meant for comparison when run on our machine whose specifications are given in C for the dataset having 10,000 objects and 500 features) during training, as per our tests using the Python package memory_profiler.
@@
We now look at other assets that DistClassiPy can offer us to help understand classification — confidence and robustness.
@@ 6.3 Feature Importance
We find that whether a feature is important or not depends not only on the classification task, but also the distance metric being used. In other words, each metric has features with which it works best, and often, these features differ for different metrics. We show in Figure 16 how the important features change for the multi-class classification task for three metrics — Canberra, Clark, and Soergel.
Figure 16: How feature importance changes for three different distance metrics (Canberra, Clark, and Soergel) in the multi-class classification. Each of these metrics has a similar performance, but there is only one feature that is common among all of them: Period_band_r — the $r$-band period.
@@ 6.4 Confidence of the distance metric classification
An interesting point to note is that, unlike the probabilistic outputs from something like RFC, our distance metric classifier with $c_{d}$ does not assume a fixed confidence level shared across all classes. The distance metric classifier does not impose a strict boundary on how far away a test object can be, and we suspect this may be useful for dealing with outliers and anomalies. However, this is not something we will explore in this work, and we end our discussion on confidences.
@@ 6.5 Robustness of the distance metric classification
We are interested in determining the robustness of our classifier — how well does our classifier perform with different types of data? We split the test set into batches, splitting the data by a given feature’s distribution into four quantiles, ensuring that each quantile contains an equal number of objects. Then, we evaluate the accuracy for each quantile separately, only using the final set of features from the SFS (section 4). Since this accuracy is now a measure of robustness, we call it the robustness score. This method is applied to all features from our “super-set” (as derived in Figure 9). For the multi-class classification task, robustness scores for the Clark and Canberra metric have been visualized as a heatmap in Figure 18. We note that, although all features in the “super-set” of the 15 best features are not used by the classifier itself, we still use them to split our quantiles, before dropping the unused feature based on the distance metric. This helps us compare the robustness for different metrics.
@@
Figure 18: Robustness scores for the multi-class classification task, with the distance metric classifier using the Clark (Figure 18a) and Canberra metrics (Figure 18b). The robustness score is calculated by splitting our test set into quantiles, and then evaluating the classification accuracy for each quantile. Note that, although each quantile is used to split the dataset, only a subset of these features are actually used in the classification (boldfaced). We still include all features from the super-set of features for the quantile split, to better compare performance for different metrics. For e.g., although the Canberra metric (b) performs better in general, Clark metric (a) performs better with objects having larger periods (quantile 4, top row). The color maps limits for Figure 18 and Figure 19 have been cut to the range [65,100] to increase the readability of differences in performance, but both models presented in this figure perform better than 78% at any quantile with any feature.
@@ 7 Conclusions
Time series classification is a non-trivial problem in machine learning. In ground-based astronomy, light curve classification is even more challenging due to the sparsity and heteroscedasticity of the data. We developed a distance metric classifier  DistClassiPy that uses distance metrics computed on features derived from light curves. Tested on multiple classification tasks, our classifier produces results comparable to other state-of-the-art feature-based methods like Random Forest Classifiers (RFCs), with $F_{1}=92\%$ and $F_{1}=94\%$ on a multi-class and a One vs. Rest classification respectively. We achieve a performance of $F_{1}=68\%$ on a particularly challenging binary classification that attempts to separate RSCVN and BYDra, two star types that exhibit photometrically similar and highly variable behaviors (Bopp, 1980), and for which, too, the RFC performance is comparable $F_{1}=70\%$. This cross-validated performance was confirmed by running the model on a “hidden” dataset completely unseen in training and that, unlike the original dataset, is not rebalanced to have an even representation of all classes. For all three classification tasks, we confirm the performance obtained in cross-validation testing: $F_{1}=93\%,92\%$, and 68% respectively for the multi-class, One vs. Rest, and binary classification.
At the same time, DistClassiPy is faster, more interpretable, and suited to be tailored to specific science goals with additional performance enhancements and computational advantages. In our tests, it performed 50×50$\times$50 × faster than RFC on small datasets (10 features by 100 objects), and over 150×150$\times$150 × faster on larger datasets (500 features by 10,000 objects) demonstrating impressive scalability properties. The dynamism in the distance metrics allows for a choice suited to the available features and the test particle’s characteristics within the feature space. Finally, because the features are a result of domain-sensitive dimensionality reduction, the user can leverage relevant knowledge in the context of the problem.